OpenStudy (zyberg):

What are all different ways to connect three kettles of 1 kW power, two of 1.5 kW and one of 2.2 kW, when maximum electrical current is 16 Amperes?

ganeshie8 (ganeshie8):

If you're living in USA, you may assume V = 120V

OpenStudy (zyberg):

Hm, perhaps we should be able to use P = VI and V = P/I? Not sure how to count how many possible ways to connect them are there tho... (And by the way, why did I learn to decode voltage as U, when everybody in other countries use V?) No, I am not living in USA :)

ganeshie8 (ganeshie8):

we use U for electric potential energy and V for electric potential

OpenStudy (zyberg):

Seems that electric potential is for one charge and electric potential energy is for overall electric charges flow?

ganeshie8 (ganeshie8):

Yes, "potential difference" is called "voltage". We use the same letter V for this too

ganeshie8 (ganeshie8):

First find out the voltage between the wires that are entering your house

OpenStudy (zyberg):

Hm, but for that I would need something more than power of kettles... or not?

ganeshie8 (ganeshie8):

I'm asking you to find out the voltage between the wires in your country

ganeshie8 (ganeshie8):

It's like a general fact, textbook authors don't bother to give this info in each and every problem...

OpenStudy (zyberg):

According to this http://www.adaptelec.com/index.php?main_page=document_general_info&products_id=161 it's 220-240 Volts

ganeshie8 (ganeshie8):

`In Lithuania the standard voltage is 230 V. The standard frequency is 50 Hz. `

ganeshie8 (ganeshie8):

Voltage = 230V use `power = (voltage)^2/(resistance)` to find the resistance of each appliance

OpenStudy (zyberg):

R = V^2/P; R_1 = 52.9; R_2 = ~35.27; R_3 = ~24.05.

ganeshie8 (ganeshie8):

Good. How many appliances do we have ?

ganeshie8 (ganeshie8):

3 of 1kW 2 of 1.5kW 1 of 2.2kW You get maximum current when you connect all these 6 appliances in parallel, right ?

OpenStudy (zyberg):

Yes.

ganeshie8 (ganeshie8):

Maybe connect them all in parallel and find out the current,

ganeshie8 (ganeshie8):

You may replace the appliances by their resistances in the ciruit diagram

OpenStudy (zyberg):

I = 3 * 230/52.9 + 2 * 230/35.27 + 230/24.05 = 35.64913335872766846251;

OpenStudy (zyberg):

It's too high.

ganeshie8 (ganeshie8):

Yes, so we cannot connect all of them in parallel. Find the current in other extreme case - connect all of them in series

OpenStudy (zyberg):

Well, then it would require each current to be equal. But when dividing to get each current we use same voltage, but different resistance, so it's not equal. How to do this?

ganeshie8 (ganeshie8):

Simplify the circuit : resistances in series add up

ganeshie8 (ganeshie8):